OpenStudy (anonymous):

Given a sphere of diameter, d. What is the percentage increase in its volume when surface area increases by 21%?

OpenStudy (perl):

LEt me start over

OpenStudy (perl):

Let d = diameter, r = radius, V for volume, S for surface area Since d = 2*r , so r = d/2 V = 4/3 Pi * r^3 = 4/3 * Pi * (d/2)^3 = Pi/ 6 * D^3 S = 4 * Pi * r^2 = 4 * pi ( d/2)^2 = pi * d^2 ok so far?

OpenStudy (perl):

the initial surface area and initial diameter we will call S1, D1. After the increase the new surface area and diameter are S2 , D2. S1 = Pi * D1^2 solve for D1 D1 = sqrt( S1 / pi ) S2 = Pi * D2^2 but we are given that S2 = 1.21*S1 therefore by substitution S2= 1.21 * Pi * D1^2 and 1.21 * Pi * D1^2=Pi * D2^2 solve for D2 D2 = sqrt( 1.21 * D1^2) Now V2 = pi/6 * D2^3. V2 = pi/6 ( sqrt(1.21*D1^2))^3 V2= pi/6 * (1.21 * D1^2)^(3/2) V2 = pi/6 * 1.21^(3/2) * D1 ^3 = (pi/6 * D1^3) * 1.21^3/2 = V1 * 1.21 ^(3/2)